OpenStudy (anonymous):

Suppose that sin theta = 3/5 and 0

OpenStudy (zehanz):

Do you know the formulas for the double angles: \(\sin 2\theta=2\sin\theta\cos\theta\) and \(\cos2\theta=2\cos^2\theta-1\) And also \(\sin^2\theta+\cos^2\theta=1\) ?

OpenStudy (anonymous):

I'm sorry I don't know.

OpenStudy (zehanz):

But you need these kind of formulas to be able to do anything with this question. I'm sure you know the last one!

OpenStudy (anonymous):

I do know the last one, however I do not know how to use them. I am barely getting to know the material.

OpenStudy (zehanz):

After all, we can't use magic, so we have to learn formulas...

OpenStudy (zehanz):

OK, starting with the last one: \(\sin^2 \theta=\frac{9}{25}\). So what is \(\cos^2\theta\) then?

OpenStudy (anonymous):

How would I know what to input for these formulas?

OpenStudy (zehanz):

You know \(\sin \theta=\frac{3}{5}\), so \(\sin^2\theta=(\sin \theta)^2=\frac{9}{25}\)

OpenStudy (anonymous):

I don't know the symbol that you put on top of the 25.

OpenStudy (zehanz):

I wrote 9/25 as \(\dfrac{9}{25}\)

OpenStudy (anonymous):

oh okay. i apologize i couldn't read that.

OpenStudy (zehanz):

No problem, I'll try to make it larger! Can you calculate with this and the third formula the value of \(\cos^2\theta\) ?

OpenStudy (anonymous):

Well I think so because you need \[\cos ^{2}\theta - \sin ^{2}\theta\] to get cos 2 theta?

OpenStudy (zehanz):

No, I think that, because \(\sin^2\theta=\dfrac{9}{25}\), that \(\cos^2\theta=\dfrac{16}{25}\), because their sum equals 1.

OpenStudy (zehanz):

This means I also have found (take the square root) \(\cos\theta=\dfrac{4}{5}\).

OpenStudy (anonymous):

What about tan theta/2?

OpenStudy (zehanz):

We cannot answer that yet, because we still haven't found the hald-angle values :(

OpenStudy (anonymous):

we would use this formula \[\tan 2 \theta = 2 \tan \theta \div 1 - \tan ^{2} \theta\]

OpenStudy (zehanz):

Remember? You were supposed to find \(\cos\dfrac{\theta}{2}\) and \(\tan\dfrac{\theta}{2}=\dfrac{\sin\dfrac{\theta}{2}}{\cos\dfrac{\theta}{2}}\). We only have the values \(\sin\theta=\dfrac{3}{5}\) and \(\cos\theta=\dfrac{4}{5}\). Found by using the last of the three formulas. So now it is time to see what the first two are for!

OpenStudy (anonymous):

Oh. Okay. So we use sin theta and cos theta to find tan theta/2?

OpenStudy (zehanz):

You might use your formula. I'm a little surprised you do know that one, and not the ones for \(\sin2\theta\) and \(\cos2\theta\)!

OpenStudy (anonymous):

Yeah I'm sorry for that. I just find math really complicated. So, how would I be able to break this down without stressing over it?

OpenStudy (zehanz):

OK, if \(\sin2\theta=2\sin\theta\cos\theta\), then \(\sin\theta=2\sin\dfrac{\theta}{2}\cos\dfrac{\theta}{2}\). Because the only thing the formula says is: angles on the left hand side are double the ones on the right hand side.

OpenStudy (zehanz):

Also, \(\cos2\theta=2\cos^2\theta-1\) means: \(\cos\theta=2\cos^2\dfrac{\theta}{2}-1\)